Televisions are advertised by the length of their diagonals. If a 42 inch measures 18 inches high, how wide is the television?
Answer by TimothyLamb(4379)

d = sqrt( ww + hh )
---
square both sides:
dd = sqrt( ww + hh )^2
dd = ww + hh
ww = dd - hh
w = sqrt( dd - hh )
---
w = sqrt( 42*42 - 18*18 )
w = 37.947331922
---
answer:
width = 37.95 inches
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---